What is the frequency of a gamma ray with λ= 2.64×10−11 m ?
1
Expert's answer
2018-06-20T05:18:49-0400
ν = c/λ, where ν is the frequency of a gamma ray, Hz; c is the speed of light, 3 x 10^8 m/s; λ is the wavelength, m

ν = 3×10^8 / 2.64×10−11 = 1.14×10^19